

********************************************************************
UARS WEEKLY STATUS REPORT - 22 MAY 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
SOLSTICE science production software v. 4.61 has been promoted to
production, and mission reprocessing of our new v. 10 data products
is underway.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and
occultation scan data without apparent problems.
***********
PEM
***********
PEM/AXIS was active from May 10 through May 15 and became active again
on May 21. It is currently generating energy deposition from inverted X-ray
spectra for each of its 16 look directions. The zenith boom instruments of PEM
we active from May 16 through May 20. While active, they generated energy
deposition profiles from in situ measured electrons and ions.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
HALOE was powered off May 19, 1998. All engineering parameters were
nominal and science channel outputs were within expected operating range.
HALOE is scheduled to be powered on June 2, 1998.
***********
MLS
***********
MLS was reactivated successfully on May 20, in the reduced 85 watt power mode.
MLS instrument operations have continued flawlessly. MLS is scheduled to
operate through June 1.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF UCSS release 10.0 has been scheduled for May 27th and should require
about 10 hours of downtime during day shift. Users were requested to send
comments, reservations, or impact statements to the PAC account.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1263C=116L)
67 days HAL (4.80) PROD (L1) (1868R-190C-2F=1676L)
1 day HAL (4.80) PROD (L3) (1868R-1C=1867L)
The CDHF plans to utilize the UCSS Scheduler enhancement (New Expansion
Tool). (PROBLEM REPORT 3427 generated 17-Jan-1997).
Currently jobs are automatically expanded every hour. This new tool will
be added to a production reprocessing DCL runstream which will speed up the
reprocessing by expanding at the completion of a job. This tool will be used
based on the requirements of the PI and how much data can be handled at the
RAC's. The CDHF will continue to use the current automatic expansion task for
all other production processing.
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 May 1998 (Orbit 36379, GMT Day 129) through 15 May
1998 (Orbit 36483, GMT Day 135).
SPACECRAFT OPERATIONS
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 129/0248 36380 On
PEM 130/0104 36394 AXIS on
WINDII 130/2339 36408 On
PEM 135/2346 36483 AXIS off
The beta angle reached a minimum angle of 0.0 degrees on 05 May 1998 (DOY
125). The beta angle is now increasing toward a maximum angle of 78.0
degrees on 26 May 1998 (DOY 146).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.1 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.1 degrees C. Additional battery performance data for
each battery on 09 May 1998 and 15 May 1998 are provided in Appendix D.
The clock error ranged from -10.2 to +10.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 04
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HALOE 02
% PEM 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/062 - 2287.496945
Transponder B: 98/062 - 2287.497010
GROUND SYSTEM OPERATION
No (0) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/127 show the observatory data loss to be
20 hours, 43 minutes, 11.156 seconds. This is a 0.0357 percent data loss
which equals a 99.9643 percent data capture for the mission.
The FOT made several minor modifications to the OBCSTF resident POCC
database this week. The FOT will spend the next several weeks testing
database 67 and making any further necessary corrections.
Identified a potential problem with the new AP software (10.11) during
flight software testing. This software was tested during the previous week
and found to support the new flight software configuration. However, the
new AP software was modified to only support the new flight software, and
is not compatible with the existing flight software configuration. The AP
software should support both configurations, so the transition to the new
flight software can be made without the need to change ground software
simultaneously. Modification of the AP software to support all flight
software configurations is being pursued.
OTHER SIGNIFICANT EVENTS
Continued testing of the upgrade to the OBC Power Monitor (PMON) flight
software. Minor changes to both the flight software and supporting ground
database were made during the week. Testing continues to proceed
successfully.
Work continues on the UARS August orbit adjust planning.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
A response to a list of questions was prepared for distribution this week.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 16 June 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 May 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- --- -------
36380 129/0248 98-158 HALOE Power ON
36394 130/0104 98-159 PEM/AXIS Power ON
36408 130/2339 98-160 WINDII Power ON
36415 131/1042 98-161 OBC Clk Adj 0.0 to 20.6 msec/day
36435 132/1846 98-163 OBC Clk Adj 20.6 to 0.0 msec/day
36440 133/0301 98-164 HALOE CMDs to reset elevation position
36469 135/0125 98-165 OBC Clk Adj 0.0 to 20.6 msec/day
36483 135/2346 98-166 PEM Power OFF AXIS
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 May 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI HRDI QB Current Monitor 36398,36404,36413
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 36379-36413
WIEUTEMP Y-LO WINDII EU Temperature 36379-36408
SOSTMPRNTRN Y-LO Temperature Entrance 36404,36480
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 May 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
36415 131/10:47 (11 May 98) from 00.0 to 20.6 msec/day
36435 132/18:58 (12 May 98) from 20.6 to 00.0 msec/day
36469 135/01:29 (15 May 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 May 1998
09 May 1998 - GMT Day 129 Beta = 15.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.94 +0.0/ +0.0 18.4 0.000 0.0
2 +100.8/ +0.0 +6.44/+3.87 +31.2/-16.0 23.8 1.020 23.2
3 +67.2/ -5.6 +5.15/+2.60 +30.4/-16.4 23.8 1.019 22.4
15 May 1998 - GMT Day 135 Beta = 40.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +56.0/ -11.2 +6.11/+4.51 +28.8/-16.0 25.1 1.023 22.5
3 +67.2/ -5.6 +5.15/+3.55 +28.0/-16.0 25.1 1.024 21.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 May 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-021 Generic Late Acq for May 1998 (Open until 31 May 1998).
SEVNTFW20